# Disentangling Representations through Multi-task Learning

### v2

@@ B.5.1 Elliptical Distributions
A multivariate distribution ([22]) is elliptical if its density function $f(\mathbf{x})$ can be expressed as:
$f(\mathbf{x})=|\boldsymbol{\Sigma}|^{-1/2}g\left((\mathbf{x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})\right)$
@@
  • Symmetric and unimodal around $\boldsymbol{\mu}$.
@@ B.5.2 Exponential Power Distributions
$f(\mathbf{x})=\frac{\beta}{2\alpha\Gamma(1/\beta)}\exp\left(-\left(\frac{|\mathbf{x}-\boldsymbol{\mu}|}{\alpha}\right)^{\beta}\right),$
where $\beta>0$ controls the kurtosis ([7]). Properties:
@@ B.5.3 Example with Multivariate t-Distribution
Suppose $\epsilon(t)$ follows a multivariate t-distribution ([39]) with degrees of freedom $\nu>2$:
@@ B.6 Discussion
Our theoretical results have important implications for the manifold hypothesis, which posits that real-world high-dimensional data tend to lie on or near low-dimensional manifolds embedded in the high-dimensional space [24, 15]. The key insight is that our proofs show an optimal multi-task classifier must encode an estimate of the disentangled coordinates of the true underlying environment state in its latent representation. Consider the disentangled space in which $\mathbf{x}^{*}$ resides, denoted $\mathcal{X}^{*}$. The injective observation map $f:\mathcal{X}^{*}\to\mathcal{X}$, where decision boundaries $y_{i}:\mathcal{X}^{*}\to\{0,1\}$ are linear. Our results imply that an optimal classifier’s latent state $\mathbf{Z}(t)$ must encode disentangled coordinates in $\mathcal{X}^{*}$ rather than ambient coordinates $\mathcal{X}$.
The injective observation map $f$ aligns closely to the typical conception of a data manifold (e.g., if $f\in C^{1}$ or $f\in C^{n}$, as described in [72]). The disentangled space $\mathcal{X}^{*}$ can be seen as the intrinsic coordinate system of the manifold, while $f$ maps these coordinates to the high-dimensional observation space $\mathcal{X}$. Our findings suggest that an optimal classifier will implicitly learn to invert this mapping to recover the disentangled coordinates. Moreover, for natural data where the manifold hypothesis holds, the learned latent representation would plausibly capture the manifold structure, as this is essential for disambiguating noisy observations and estimating the true underlying state. The low-dimensional manifold structure is a key prior that an optimal classifier can (and in our case must) exploit to improve its performance.
@@
Consider an analogy with masked autoencoder vision foundation models, where $\mathbf{x}^{*}$ is the “ground truth” of a scene (objects, positions, states, and relationships), the measurement variable $\mathbf{X}$ is an image with missing patches [19, 29], and the model predicts the missing patch data $y(\mathbf{x}^{*})$. The model’s latent variable $\mathbf{Z}$ exhibits some “understanding” of $\mathbf{x}^{*}$ in the form of abstract representations useful for downstream tasks. This analogy extends to masked language models [17] and autoregressive language models [62], where $\mathbf{x}^{*}$ is “meaning” in a semantic space, $\mathbf{X}(t)$ are words, and $y$ is the next word. Localizing $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$ relates to constructing a world model, showing that $\mathbf{Z}$ represents $\mathbf{x}^{*}$ abstractly and with high fidelity.
@@
The curvature of the observation map $f$ and the level of noise $\sigma$ are fundamental factors influencing the extent of the degeneracy introduced by the noise after the observation map. High curvature in $f$ can make the intrinsic geometry of the data more challenging to identify (e.g., more tightly crumpled paper). Large noise levels can push observations further from the underlying manifold, similarly worsening the potential degeneracy in the observations. The reach of the manifold $f$ ([1]) can be used as an immediate loose bound for post-observation map noise $\epsilon_{2}(t)$ to ensure that the derived theorems still hold. Interestingly, this observation map degeneracy is a key concern in work on visual predictive coding for map learning by Gornet and Thomson [27]. They demonstrate that a predictive coding network can learn to map a Minecraft environment with visually degenerate states by integrating information to perform predictive coding along trajectories within the environment.
@@
Our results provide a new perspective on the Platonic representation hypothesis [34]. The Platonic representation hypothesis suggests that the convergence in deep neural network representations is driven by a shared statistical model of reality, like Plato’s concept of an ideal reality. Convergence of representations is analyzed in terms of similarity of distances between embedded datapoints among AI models trained on various modalities. While the authors of the hypothesis argue that energy constraints might lead to divergence from a shared representation for specialized tasks, our Optimal Representation Theorem suggests that the key factor driving convergence is the diversity and comprehensiveness of the tasks being learned. As long as the tasks collectively span the space of the underlying data representation, convergence to a shared, reality-aligned representation can occur, even in the presence of energy or computational limitations. Our theoretical results amount to a necessary condition for optimal multi-task classifiers to represent a disentangled representation of the data within their latent state. With energy constraints, extraneous network activity may be regularized out of the model, resulting in greater alignment between disentangled representations in energy constrained models that “understand” the Platonic nature of reality. The very energy constraints [34] suggest may lead to divergence could actually facilitate convergence of the platonic representations, as they may encourage the learning of simple, generalizable features that capture the essential structure of the data. This insight opens up interesting avenues for future research on the interplay between task diversity, energy constraints, and the emergence of shared representations. Finally, energy constraints have been shown to naturally lead to predictive coding [2], tightening the relationship between energy efficiency, prediction, and cognitive map learning.

### v3

@@ B.5.1 Elliptical Distributions
A multivariate distribution (Fang (2018)) is elliptical if its density function $f(\mathbf{x})$ can be expressed as:
$f(\mathbf{x})=|\bm{\Sigma}|^{-1/2}g\left((\mathbf{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right)$
@@
  • Symmetric and unimodal around $\bm{\mu}$.
@@ B.5.2 Exponential Power Distributions
$f(\mathbf{x})=\frac{\beta}{2\alpha\Gamma(1/\beta)}\exp\left(-\left(\frac{|\mathbf{x}-\bm{\mu}|}{\alpha}\right)^{\beta}\right),$
where $\beta>0$ controls the kurtosis (Box & Tiao (2011)). Properties:
@@ B.5.3 Example with Multivariate t-Distribution
Suppose $\epsilon(t)$ follows a multivariate t-distribution (Kotz & Nadarajah (2004)) with degrees of freedom $\nu>2$:
@@ B.6 Discussion
Our theoretical results have important implications for the manifold hypothesis, which posits that real-world high-dimensional data tend to lie on or near low-dimensional manifolds embedded in the high-dimensional space (Fefferman et al., 2016; Olah, 2014). The key insight is that our proofs show an optimal multi-task classifier must encode an estimate of the disentangled coordinates of the true underlying environment state in its latent representation. Consider the disentangled space in which $\mathbf{x}^{*}$ resides, denoted $\mathcal{X}^{*}$. The injective observation map $f:\mathcal{X}^{*}\to\mathcal{X}$, where decision boundaries $y_{i}:\mathcal{X}^{*}\to\{0,1\}$ are linear. Our results imply that an optimal classifier’s latent state $\mathbf{Z}(t)$ must encode disentangled coordinates in $\mathcal{X}^{*}$ rather than ambient coordinates $\mathcal{X}$.
The injective observation map $f$ aligns closely to the typical conception of a data manifold (e.g., if $f\in C^{1}$ or $f\in C^{n}$, as described in Tu (2017)). The disentangled space $\mathcal{X}^{*}$ can be seen as the intrinsic coordinate system of the manifold, while $f$ maps these coordinates to the high-dimensional observation space $\mathcal{X}$. Our findings suggest that an optimal classifier will implicitly learn to invert this mapping to recover the disentangled coordinates. Moreover, for natural data where the manifold hypothesis holds, the learned latent representation would plausibly capture the manifold structure, as this is essential for disambiguating noisy observations and estimating the true underlying state. The low-dimensional manifold structure is a key prior that an optimal classifier can (and in our case must) exploit to improve its performance.
@@
Consider an analogy with masked autoencoder vision foundation models, where $\mathbf{x}^{*}$ is the “ground truth” of a scene (objects, positions, states, and relationships), the measurement variable $\mathbf{X}$ is an image with missing patches (Dosovitskiy et al., 2020; He et al., 2021), and the model predicts the missing patch data $y(\mathbf{x}^{*})$. The model’s latent variable $\mathbf{Z}$ exhibits some “understanding” of $\mathbf{x}^{*}$ in the form of abstract representations useful for downstream tasks. This analogy extends to masked language models (Devlin et al., 2018) and autoregressive language models (Radford et al., 2019), where $\mathbf{x}^{*}$ is “meaning” in a semantic space, $\mathbf{X}(t)$ are words, and $y$ is the next word. Localizing $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$ relates to constructing a world model, showing that $\mathbf{Z}$ represents $\mathbf{x}^{*}$ abstractly and with high fidelity.
@@
The curvature of the observation map $f$ and the level of noise $\sigma$ are fundamental factors influencing the extent of the degeneracy introduced by the noise after the observation map. High curvature in $f$ can make the intrinsic geometry of the data more challenging to identify (e.g., more tightly crumpled paper). Large noise levels can push observations further from the underlying manifold, similarly worsening the potential degeneracy in the observations. The reach of the manifold $f$ (Aamari et al. (2019)) can be used as an immediate loose bound for post-observation map noise $\epsilon_{2}(t)$ to ensure that the derived theorems still hold. Interestingly, this observation map degeneracy is a key concern in work on visual predictive coding for map learning by Gornet & Thomson (2024). They demonstrate that a predictive coding network can learn to map a Minecraft environment with visually degenerate states by integrating information to perform predictive coding along trajectories within the environment.
@@
Our results provide a new perspective on the Platonic representation hypothesis (Huh et al., 2024). The Platonic representation hypothesis suggests that the convergence in deep neural network representations is driven by a shared statistical model of reality, like Plato’s concept of an ideal reality. Convergence of representations is analyzed in terms of similarity of distances between embedded datapoints among AI models trained on various modalities. While the authors of the hypothesis argue that energy constraints might lead to divergence from a shared representation for specialized tasks, our Optimal Representation Theorem suggests that the key factor driving convergence is the diversity and comprehensiveness of the tasks being learned. As long as the tasks collectively span the space of the underlying data representation, convergence to a shared, reality-aligned representation can occur, even in the presence of energy or computational limitations. Our theoretical results amount to a necessary condition for optimal multi-task classifiers to represent a disentangled representation of the data within their latent state. With energy constraints, extraneous network activity may be regularized out of the model, resulting in greater alignment between disentangled representations in energy constrained models that “understand” the Platonic nature of reality. The very energy constraints Huh et al. (2024) suggest may lead to divergence could actually facilitate convergence of the platonic representations, as they may encourage the learning of simple, generalizable features that capture the essential structure of the data. This insight opens up interesting avenues for future research on the interplay between task diversity, energy constraints, and the emergence of shared representations. Finally, energy constraints have been shown to naturally lead to predictive coding Rao & Ballard (1999); Ali et al. (2022), tightening the relationship between energy efficiency, prediction, and cognitive map learning. A relationship between predictive coding and optimal Bayesian estimation has also been established (Rao, 1999).